the height of this oversized container is 3m. its diameter is 2m. what is its area?

Answer by dabanfield(803)

The surface area would be the circumference times the height (C*H).
The circumference C is given by pi*D where D is the diameter so:
C=pi*2
The height is 3 so the surface area is:
C*H = (pi*2)*3 = 6*pi